OpenStudy (anonymous):

HELP WANTED!!! (sin 24°) (csc 24°)

OpenStudy (cwrw238):

csc = 1 / sin

OpenStudy (australopithecus):

since csc = 1/sin(x) we have sin(24) * (1/sin(24))

OpenStudy (australopithecus):

you dont need a calculator to solve this

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

THE CSC HAD ME CONFUSED

OpenStudy (australopithecus):

Do you know SOH CAH TOA?

OpenStudy (cwrw238):

welcome

OpenStudy (australopithecus):

just remember the order Sin, Cos, Tan CSC Sec Cot I just do this to remember SCT CSC

OpenStudy (australopithecus):

CSC = 1/sin Cos = 1/cos Cot = 1/tan or cos/sin

OpenStudy (anonymous):

thAt HELPS

OpenStudy (australopithecus):

Always look for patterns and try to link stuff it helps a lot in math and school in general

OpenStudy (anonymous):

OK I WILL